# Infer-and-widen versus split-and-condition: two tales of selective inference

### v1

@@ 3 Vignette #1: The winner’s curse
Figure 1: Vignette #1, under the model (3.1) with $\mu=0$ and $\sigma=1$ and selection rule in 3.3. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of widths of the infer-and-widen 3.4 and split-and-condition 3.6 intervals with $95\%$ coverage across values of the sample size $n$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $n=100$, we display the width of the “oracle” infer-and-widen interval and the split-and-condition interval 3.6 as a function of coverage. The 95% classical interval, and the “attainable” 95% infer-and-widen interval based on algorithmic stability 3.4, are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
@@
$\mathcal{S}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}(Y_{i}+\zeta_{i}),\;\;\;\;\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right).$ (3.3)
@@
Here, $c$ is a positive constant that governs the amount of additive noise. When $c=0$, $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ simplifies to $\mathcal{S}(Y)$ in (3.2).
@@
Proposition 1 (An infer-and-widen interval based on (3.3), motivated by algorithmic stability (Zrnic and Jordan, 2023)).
@@
Under the model (3.1), consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(\cdot)$ defined in (3.3). For any $\eta>0$ and $\nu\in(0,\alpha)$ such that $\eta c\geq 2z_{1-\alpha(\alpha-\nu)/2n}$, where $z_{q}$ denotes the $q$th quantile of the standard normal distribution, a valid infer-and-widen $1-\alpha$ confidence interval for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ takes the form
$\mathrm{CI}_{{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\alpha+\nu)e^{-\eta}}(Y):=\left[Y_{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}\pm\sigma z_{1-\alpha(1-\alpha+\nu)e^{-\eta}/2}\right].$ (3.4)
We will now construct a split-and-condition interval that makes use of exactly the same selection event, (3.3). Towards this end, note that $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)=\mathcal{S}(Y^{\mathrm{train}})$ where $Y^{\mathrm{train}}:=Y+\zeta$.
@@
Since $Y\mid Y^{\mathrm{train}}$ does not admit a tractable distribution, we will additionally condition on ${\rm sign}(\zeta_{1}),\ldots,{\rm sign}(\zeta_{n})$, as in (2.3). This yields a multivariate truncated normal distribution with independent marginals, which we use to obtain the confidence interval in Proposition 2.
Proposition 2 (A split-and-condition interval based on (3.3)).
@@
Under the model (3.1), consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ defined in (3.3). Define $\delta_{i}:={\rm sign}(\zeta_{i})$ and $A_{i}:=\{y_{i}\in\mathbb{R}:\delta_{i}y_{i}\leq\delta_{i}(Y_{i}+\zeta_{i})\}$. Let $\Phi_{\mu,\sigma^{2},A_{i}}(\cdot)$ denotes the CDF of a $N(\mu,\sigma^{2})$ distribution, truncated to the set $A_{i}$, and define $a_{i}$ and $b_{i}$ to satisfy
$\Phi_{a_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=1-\alpha/2,\quad\text{and}\quad\Phi_{b_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=\alpha/2.$ (3.5)
A valid split-and-condition $1-\alpha$ confidence interval for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ takes the form
@@
$\mathrm{CI}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}(Y_{i}\mid Y+\zeta,\delta):=\left[a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)},b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right].$ (3.6)
In the notation of Table 1, (3.6) uses $\mathcal{C}(Y)=(Y^{\mathrm{train}},\delta)$. This procedure lies somewhere between data fission (Leiner et al., 2023) and randomized conditional selective inference (Tian and Taylor, 2018; Panigrahi et al., 2024) in Table 1, in that we conduct selection using $Y^{\mathrm{train}}$, and then condition on a bit more for inference. Computing the confidence interval in Proposition 2 requires numerically solving (3.5) for $a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ and $b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$.
@@
Thus, both (3.4) and (3.6) provide confidence intervals for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that attain coverage $1-\alpha$. We should prefer the approach that yields narrower intervals. Figure 1(a) displays the ratio of the width of the infer-and-widen confidence interval (3.4) to the width of the split-and-condition confidence interval (3.6) when $\mu=0$ and $\sigma^{2}=1$, as we vary the sample size $n$ and the Laplacian variance $2c^{2}$ in (3.3).
@@
The infer-and-widen interval (3.4) involves tuning parameters $\nu$ and $\eta$; to construct Figure 1(a), for a fixed value of the noise $c$ and sample size $n$ in (3.3), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $n$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (3.4) of apparently infinite width. The split-and-condition intervals are much narrower than the infer-and-widen intervals across all values of $n$ and $c$ considered.
@@
It is possible that the width of the infer-and-widen interval (3.4) is due to the fact that Zrnic and Jordan (2023) apply a series of inequalities in its derivation. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $Y_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value $\mu$. In Figures 1(b, c), we numerically compare the “oracle” infer-and-widen method to the split-and-condition interval (3.6) under the selection rule (3.3) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is substantially wider than the split-and-condition interval in (3.6), thereby indicating that no infer-and-widen interval can outperform the split-and-condition interval (3.4) in this setting.
@@ 4 Vignette #2: Maximal contrasts
$\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X^{\top}_{j}Y+\zeta_{j}\right|,\quad\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}\mathrm{Laplace}\left(c\right),$ (4.2)
where $c$ is a positive constant that governs the scale of additive noise.
Proposition 1 (An infer-and-widen interval based on (4.2), motivated by algorithmic stability (Zrnic and Jordan, 2023)).
Under the model (3.1), consider the Laplacian selection rule $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ defined in (4.2). For any $\eta>0$ and $\nu\in(0,1)$ such that $c\geq 2z_{1-\alpha\nu/(2p)}/\eta$, a valid infer-and-widen $1-\alpha$ confidence interval for $X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\mu$ takes the form
$\mathrm{CI}_{{{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\nu)e^{-\eta}}(Y):=\left[X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y\pm\sigma z_{1-\alpha(1-\nu)e^{-\eta}/2}\right].$ (4.3)
We now construct a split-and-condition interval that makes use of the selection event (4.2). Note that $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ is a function of $X^{\top}Y+\zeta$. Since the distribution of $X^{\top}Y\mid X^{\top}Y+\zeta$ is intractable, we cannot directly use it to derive a split-and-condition interval, as in Proposition 2. Proposition 2 instead makes use of the distribution of $X^{\top}Y$ conditional on $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ and some additional quantities.
Proposition 2 (A split-and-condition interval based on 4.2).
@@
$W$ $:=(I_{n}-X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top})Y$
@@
$\delta$ $:={\rm sign}\left(X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)$
$v_{\max}$ $:=\underset{j:\left(\delta AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}<0}{\mathrm{max}}-(A\zeta+AX^{\top}W)_{j}/(AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)})_{j}$
$v_{\min}$ $:=\min_{j:\left(\delta AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}>0}-(A\zeta+AX^{\top}W)_{j}/(AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)})_{j},$
where $A\in\mathbb{R}^{(2p+1)\times p}$ is defined in Subsection B.1. Let $\Phi_{\lambda,\sigma^{2},[v_{\min},v_{\max}]}(\cdot)$ denote the CDF of the normal distribution with mean $\lambda$ and variance $\sigma^{2}$, truncated to the set $[v_{\min},v_{\max}]$, and define $a_{j}$ and $b_{j}$ to satisfy
$\Phi_{a_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=1-\alpha/2\quad\text{and}\quad\Phi_{b_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=\alpha/2.$ (4.4)
A $1-\alpha$ split-and-condition confidence interval for $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$ takes the form
@@
$\mathrm{CI}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y\mid\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta\right):=\left[a_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)},b_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right].$ (4.5)
@@
In the notation of Table 1, (4.5) uses $\mathcal{C}(Y)=(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta)$. Computing 4.5 requires numerically solving 4.4 for $a_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}$ and $b_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}$.
@@
Figure 2: Vignette #2, under the model (3.1) with $n=100$, $\sigma=1$, $\mu=X^{\top}\phi$, and selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of widths of infer-and-widen 4.3 and split-and-condition 4.5 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $p=100$, we display the width of the “oracle” infer-and-widen interval and the split-and-condition interval 4.5 as a function of coverage. The 95% classical interval, and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability, are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
@@
We now compare the interval (4.3) to (4.5), as well as to the “oracle” infer-and-widen interval using the Laplace selection rule (4.2). We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X^{\top}\phi$ where half of the elements of $\phi$ are zero and the other half are drawn from an exponential distribution with mean $50/7$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and the columns of $X$ are scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 2(a) displays the ratio of confidence interval widths across values of $p$ and the variance of the Laplacian noise $2c^{2}$ in 4.2.
@@
The infer-and-widen interval (4.5) involves tuning parameters $\nu$ and $\eta$; to construct Figure 2(a), for a fixed value of the noise $c$ and number of features $p$ in (4.2), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $p$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (4.3) of apparently infinite width. The split-and-condition intervals are narrower than the infer-and-widen intervals across most values of $p$ and $c$ considered.
@@
It is possible that the width of the infer-and-widen interval (4.3) is due to the fact that Zrnic and Jordan (2023) apply a series of inequalities in its derivation. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value. In Figures 2(b, c), we numerically compare the “oracle” infer-and-widen method to the split-and-condition interval (4.5) under the selection rule (4.2) with variance $0.332c^{2}=$ 0.33 and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is narrower than the split-and-condition interval in (4.5), thereby indicating that in principle, an infer-and-widen interval can outperform the split-and-condition interval (4.5) in this setting. However, to the best of our knowledge, no methods to construct such an infer-and-widen interval are currently available.
Remark 1.
Figure 2 of Zrnic and Jordan (2023) shows that the infer-and-widen interval (4.3) far outperforms a split-and-condition interval based on sample splitting in Vignette #2. In fact, in this setting, a split-and-condition interval based on sample splitting is pathological and better split-and-conditon intervals are available; see Subsection B.2. Thus, our results are not at odds with those of Zrnic and Jordan (2023).

### v2

@@ 3 Vignette #1: The winner’s curse
Figure 1: Vignette #1 under (3.1) with $\mu=0$, $\sigma=1$, and the selection rule in 3.3. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of the infer-and-widen 3.4 and data fission (DF) 3.7 intervals with $95\%$ coverage across values of the sample size $n$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $n=100$, we display the average width of the “oracle” infer-and-widen interval and the data fission interval 3.7 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval based on algorithmic stability 3.4 are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
@@
$\mathcal{S}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}(Y_{i}+\zeta_{i}),$ (3.3)
@@
where $\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$ and $c$ is a positive constant that governs the amount of additive noise. When $c=0$, $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ simplifies to $\mathcal{S}(Y)$ in (3.2).
@@
Proposition 1 (An infer-and-widen interval based on (3.3), given by (Zrnic and Jordan, 2023)).
@@
Under 3.1, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(\cdot)$ defined in (3.3). For any $\eta>0$ and $\nu\in(0,\alpha)$ such that $\eta c\geq 2z_{1-\alpha(\alpha-\nu)/2n}$, where $z_{q}$ denotes the $q$th quantile of the standard normal distribution, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\alpha+\nu)e^{-\eta}}(Y):=\left[Y_{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}\pm\sigma z_{1-\alpha(1-\alpha+\nu)e^{-\eta}/2}\right]$ (3.4)
has $1-\alpha$ unconditional selective coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 2.1.
@@
We will now construct an alternative interval that makes use of exactly the same selection event, (3.3), using ideas from the data fission proposal of Leiner et al. (2023). Towards this end, note that $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)=\mathcal{S}(Y^{\mathrm{train}})$ where $Y^{\mathrm{train}}:=Y+\zeta$. Since $Y\mid Y^{\mathrm{train}}$ does not admit a tractable distribution, we will additionally condition on $\left({\rm sign}(\zeta_{1}),\ldots,{\rm sign}(\zeta_{n})\right)$. This yields a multivariate truncated normal distribution with independent marginals.
Proposition 2 (A data fission interval based on (3.3)).
@@
Under (3.1), consider the selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ defined in (3.3). Define $\delta_{i}:={\rm sign}(\zeta_{i})$ and $A_{i}:=\{y_{i}\in\mathbb{R}:\delta_{i}y_{i}\leq\delta_{i}(Y_{i}+\zeta_{i})\}$. Let $\Phi_{\mu,\sigma^{2},A_{i}}(\cdot)$ denote the CDF of a ${\bf{\rm N}}\,(\mu,\sigma^{2})$ distribution truncated to the set $A_{i}$, and define $a_{i}$ and $b_{i}$ to satisfy
$\Phi_{a_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=1-\alpha/2$ (3.5)
and
$\Phi_{b_{i}+\delta_{i}\sigma^{2}/c,\sigma^{2},A_{i}}(Y)=\alpha/2.$ (3.6)
The confidence interval
$\mathrm{CI}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}(Y):=\left[a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)},b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (3.7)
@@
has $1-\alpha$ conditional selective coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(Y+\zeta,\delta\right).$
Computing (3.7) requires numerically solving (3.5) for $a_{i}$ and (3.6) for $b_{i}$, where $i=\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$.
@@
Thus, both (3.4) and (3.7) provide confidence intervals for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that attain $1-\alpha$ unconditional selective coverage. We should prefer the approach that yields narrower intervals. Figure 1(a) displays the ratio of the widths of the two intervals when $\mu=0$ and $\sigma^{2}=1$, as we vary the sample size $n$ and the Laplacian variance $2c^{2}$ in (3.3).
@@
The infer-and-widen interval (3.4) involves tuning parameters $\nu$ and $\eta$; to construct Figure 1(a), for a fixed value of the noise $c$ and sample size $n$ in (3.3), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $n$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (3.4) of apparently infinite width. The data fission intervals (3.7) are much narrower than the infer-and-widen intervals (3.4), across all values of $n$ and $c$ considered.
@@
It is possible that the width of the infer-and-widen interval (3.4) is due to the fact that Zrnic and Jordan (2023) apply a series of inequalities in its derivation. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $Y_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value $\mu$. In Figures 1(b, c), we numerically compare the “oracle” infer-and-widen method to the data fission interval (3.7) under the selection rule (3.3) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. The former method has a wider average width than the latter: thus, no infer-and-widen interval can outperform the data fission interval (3.4) in this setting.
@@ 4 Vignette #2: Maximal contrasts
$\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X^{\top}_{j}Y+\zeta_{j}\right|,$ (4.2)
where $\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$ and $c$ is a positive constant that governs the scale of the additive noise.
Proposition 1 (An infer-and-widen interval based on (4.2), given by Zrnic and Jordan (2023)).
Under (3.1), consider the Laplacian selection rule $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ defined in (4.2). For any $\eta>0$ and $\nu\in(0,1)$ such that $c\geq 2z_{1-\alpha\nu/(2p)}/\eta$, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\nu)e^{-\eta}}(Y):=\left[X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y\pm\sigma z_{1-\alpha(1-\nu)e^{-\eta}/2}\right]$ (4.3)
has $1-\alpha$ unconditional selective coverage for $X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\mu$ in the sense of 2.1.
We will take a randomized conditional selective inference approach (Tian and Taylor, 2018; Panigrahi et al., 2024) to construct an alternative confidence interval that makes use of the identical selection event.
Proposition 2 (A randomized conditional selective inference interval based on 4.2).
@@
$W$ $:=(I_{n}-X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top})Y,$
@@
$\delta$ $:={\rm sign}\left(X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right),$
and $a_{j}$ and $b_{j}$ to satisfy
$\Phi_{a_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=1-\alpha/2$ (4.4)
and
$\Phi_{b_{j},\sigma^{2},[v_{\min},v_{\max}]}(X_{j}^{\top}Y)=\alpha/2,$ (4.5)
where $\Phi_{\lambda,\sigma^{2},[v_{\min},v_{\max}]}(\cdot)$ is defined in Proposition 2, and $v_{\min}$ and $v_{\max}$ are defined in LABEL:sec:deferred. The confidence interval
$\mathrm{CI}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}\left(Y\right):=\left[a_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)},b_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (4.6)
@@
has $1-\alpha$ conditional selective coverage for $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta\right)$.
@@
Computing 4.6 requires numerically solving 4.4 for $a_{i}$ and 4.5 for $b_{i}$, where $i=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$.
@@
Figure 2: Vignette #2 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of infer-and-widen 4.3 and randomized conditional selective inference (RCSI) 4.6 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $p=100$, we display the average width of the “oracle” infer-and-widen interval and the RCSI interval 4.6 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
@@
We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$ where half of the elements of $\phi$ are zero and the other half are drawn from an exponential distribution with mean $50/7$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and its columns are subsequently scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 2(a) displays the ratio of confidence interval widths across values of $p$ and the variance of the Laplacian noise $2c^{2}$ in 4.2.
@@
The infer-and-widen interval (4.6) involves tuning parameters $\nu$ and $\eta$; to construct Figure 2(a), for a fixed value of the noise $c$ and number of features $p$ in (4.2), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $p$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (4.3) of infinite width. The randomized conditional selective inference intervals are narrower than the infer-and-widen intervals across most values of $p$ and $c$ considered.
@@
Once again, we consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y$ that achieves a given coverage. In Figures 2(b, c), we numerically compare the “oracle” infer-and-widen method to the randomized conditional selective inference interval (4.6) under the selection rule (4.2) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is narrower than the randomized conditional selective inference interval in (4.6), thereby indicating that in principle, an infer-and-widen interval can outperform the randomized conditional selective inference interval (4.6) in this setting. However, to the best of our knowledge, no methods to construct such a narrow infer-and-widen interval are currently available.